A list of numbers consists of p positive and n negative numbers. If a numbers is picked at random from this list, the probability that the number is positive is 3/5. What is the value of n/p?
A. 3/8 B. 5/8 C. 2/3 D. 3/2 E. 8/3

Answer by london maths tutor(243)


3/5 means there are 3 positive numbers out of 5 numbers
Therefore, 5-3 = 2 (negative numbers)
n = 2 and p = 3
Therefore, n/p = 2/3
Answer: (c)